OpenStudy (anonymous):

Two runners are saving money to attend a marathon. The first runner has $110 in savings, received a $45 gift from a friend, and will save $25 each month. The second runner has $50 in savings and will save $60 each month. After how many months will both runners have the same amount of money? 2 3 4 5

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

3 months

OpenStudy (anonymous):

155+25+25+25=230 50+60+6+60=230

OpenStudy (anonymous):

110 + 45 + 25 m =50 + 60 m m= 3 months